OpenStudy (theresa.womack):

How would you solve √2x-3 +5=6?

OpenStudy (anonymous):

\[ \sqrt{2}x - 3 + 5 = 6\] \[\sqrt{2}x = 6 - 5 + 3\] \[x = 4\div \sqrt{2}\] \[x=4.\sqrt{2}/\sqrt{2}.\sqrt{2}\] \[x = 2\sqrt{2}\]